A buffer solution is prepared by mixing 30.0 mL of 0.132 mol L-1 lactic acid (pKa = 3.86) and 70.0 mL of 0.243 mol L-1 sodium lactate. What is the pH of the buffer solution?
1
Expert's answer
2019-05-13T04:59:22-0400

We will use this formula to count out the pH of the buffer solution.

pH = pKa + lg(c(C3H5NaO3) * V(C3H5NaO3) / c(C3H6O3) * V(C3H6O3)) = 3.86 + lg(0.234 mol/L * 0.07 L / 0.123 mol/L * 0.03 L) = 3.86 + lg(0.01701/0.00396) = 3.86 + lg4.295 = 3.86 + 0.633 = 4.493.